OpenStudy (anonymous):

Problem: Rosalie is organizing a circus performance to raise money for a charity. She is trying to decide how much to charge for tickets. From past experience, she knows that the number of people who will attend is a linear function of the price per ticket. If she charges $5, 1200 people will attend. If she charges $7, 970 people will attend. How much should she charge per ticket to make the most money?

OpenStudy (anonymous):

If you consider these numbers as points, where x = ticket price and y = people then you have P1 = (4, 1200) and P2 = (7, 970) Determine the equation of a line that passes through these points. You will obtain a linear equation of the form y = mx + b. Notice, though, that Profit itself is a function of x and y. In fact Profit = x * y (it equals the number of people multiplied by the ticket price). So your profit equation would be P(x) = x * y = x * (mx + b) Determine the vertex of this profit equation. This will be the point, x (the ticket price) that results in the maximum profit P(x).

OpenStudy (anonymous):

What is the exact equation?

OpenStudy (anonymous):

I meant to say P1 = (5, 1200) ******** Well if the points are (5, 1200) and (7, 970), the slope is (1200 - 970) / (5 - 7) = -115 In point slope form we can write y - y1 = m(x - x1) y - 970 = -115(x - 7) y = -115x + 1775 Our profit is x * y or x * (-115x + 1775) Profit = P(x) = -115x^2 + 1775x This is a parabola with a maximum value of P(x) where x is the x-coordinate of the vertex. Can you determine the vertex of this parabola?